Give some typical numbers for tire pressures generally used in Boeing 707 and Boeing 747 planes. For their main assembly calculate and draw the tire imprint area for circular and rectangular shapes.

Type of Plane   

Tire Pressure (psi)

Boeing 707-320C   

180

Boeing 707-120B

170

Boeing 747

204

Boeing 707-320C, Maximum load on each main assembly = 157000 lb (4 Tire)

Boeing 707-120B, Maximum load on each main assembly = 120000 lb (4 Tire)

Boeing 747, Maximum load on each main assembly = 166500 lb (8 Tire)
Answer #1

The tyre pressure generally ued in boeing 707 and 747 is 200 psi.

Calculation for Boeing 707 320C :

Load on each tyre= 157000/4 = 39250 lbs

Area of imprint of each tyre = 39250/180 = 218.056 in2

Circular Radius = (Area/3.14)0.5 = 8.33 in

Square side = (Area)0.5= 14.77 in

Calculation for Boeing 707 120B :

Load on each tyre = 120000/4 = 30000 lbs

Area of imprint of each tyre = 30000/170 = 176.47 in2

Circular Radius =(Area/3.14)0.5 = 7.495 in

Square Side = (Area)0.5 = 13.28 in

Calculation for Boeing 747 :

Load on each tyre = 166500/8 = 20812.5 lbs

Area of imprint of each tyre = 20812.5/204 = 102.022 in2​​​​​​​

Circular Radius = (Area/3.14)0.5 = 5.7 in

Square Side =  (Area)0.5 = 10.1 in